Sale plants 4 saplings, in a row, in her garden. The distance between two adjacent saplings is ¾ m. Find the distance between the first and the last sapling.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question, it is given that,
The distance between two adjacent saplings = 3⁄4 m
Number of saplings planted by Saili in a row = 4
Then, number of gap in saplings = 3⁄4 × 4 =3
∴The distance between the first and the last saplings = 3 × 3⁄4
= (9/4) m
= 2 whole 1⁄4 m
Hence, the distance between the first and the last saplings is 2 whole 1⁄4 m.

Thanks